# Private Heterogeneous Federated Learning Without a Trusted Server Revisited: Error-Optimal and Communication-Efficient Algorithms for Convex Losses

### v1

@@ Abstract
We revisit the problem of federated learning (FL) with private data from people who do not trust the server or other silos/clients. In this context, every silo (e.g. hospital) has data from several people (e.g. patients) and needs to protect the privacy of each person’s data (e.g. health records), even if the server and/or other silos try to uncover this data. Inter-Silo Record-Level Differential Privacy (ISRL-DP) prevents each silo’s data from being leaked, by requiring that silo $i$’s communications satisfy item-level differential privacy. Prior work Lowy and Razaviyayn (2023a) characterized the optimal excess risk bounds for ISRL-DP algorithms with homogeneous (i.i.d.) silo data and convex loss functions. However, two important questions were left open: (1) Can the same excess risk bounds be achieved with heterogeneous (non-i.i.d.) silo data? (2) Can the optimal risk bounds be achieved with fewer communication rounds? In this paper, we give positive answers to both questions. We provide novel ISRL-DP FL algorithms that achieve the optimal excess risk bounds in the presence of heterogeneous silo data. Moreover, our algorithms are more communication-efficient than the prior state-of-the-art. For smooth loss functions, our algorithm achieves the optimal excess risk bound and has communication complexity that matches the non-private lower bound. Additionally, our algorithms are more computationally efficient than the previous state-of-the-art.
@@ 1 Introduction
Federated learning (FL) is a distributed machine learning paradigm in which multiple silos (a.k.a. clients), such as hospitals or cell-phone users, collaborate to train a global model. In FL, silos store their data locally and exchange focused updates (e.g. stochastic gradients), sometimes making use of a central server (Kairouz et al., 2021). FL has been applied in myriad domains, from consumer digital products (e.g. Google’s mobile keyboard Hard et al. (2018) and Apple’s iOS Apple (2019)) to healthcare Courtiol et al. (2019), finance FedAI (2019), and large language models (LLMs) Hilmkil et al. (2021).
@@
Differential privacy (DP) Dwork et al. (2006) guarantees that private data cannot be leaked. Different variations of DP have been considered for FL. Central DP prevents the final trained FL model from leaking data to an external adversary Jayaraman and Wang (2018); Noble et al. (2022). However, central DP has two major drawbacks: (1) it does not provide a privacy guarantee for each individual silo; and (2) it does not ensure privacy when an attacker/eavesdropper has access to the server or to another silo.
Another notion of DP for FL is user-level DP McMahan et al. (2018); Geyer et al. (2017); Levy et al. (2021). User-level DP mitigates the drawback (1) of central DP, by providing privacy for the complete local data set of each silo. User-level DP is practical in the cross-device FL setting, where each silo is a single person (e.g. cell-phone user) with a large number of records (e.g. text messages). However, user-level DP still permits privacy breaches if an adversary has access to the server or eavesdrops on the communications between silos. Moreover, user-level DP is not well-suited for cross-silo federated learning, where silos represent organizations like hospitals, banks, or schools that house data from many individuals (e.g. patients, customers, or students). In the cross-silo FL context, each person possesses a record, referred to as an “item,” which may include sensitive data. Therefore, an appropriate notion of DP for cross-silo FL should safeguard the privacy of each individual record (i.e. “item-level differential privacy”) within silo $i$, rather than the complete aggregated data of silo $i$.
Following Lowy and Razaviyayn (2023a); Lowy et al. (2023a); Liu et al. (2022) (among others), this work considers inter-silo record-level DP (ISRL-DP). ISRL-DP requires that the full transcript of messages sent by silo $i$ satisfy item-level DP, for all silos $i$. Thus, ISRL-DP guarantees the privacy of each silo’s local data, even in the presence of an adversary with access to the server, other silos, or the communication channels between silos and the server. See Figure 1 for an illustration. If each silo only has one record, then ISRL-DP reduces to local DP Kasiviswanathan et al. (2011); Duchi et al. (2013). However, in the FL setting where each silo has many records, ISRL-DP FL is more appropriate and permits much higher accuracy than local DP Lowy and Razaviyayn (2023a). Moreover, ISRL-DP implies user-level DP if the ISRL-DP parameters are sufficiently small Lowy and Razaviyayn (2023a). Thus, ISRL-DP is a practical privacy notion for cross-silo and cross-device FL when individuals do not trust the server or other silos/clients/devices with their sensitive data.
@@
Assume that the samples $\{x_{i,j}\}_{i\in[N],j\in[n]}$ are independent. The FL problem is homogeneous if the data is i.i.d., i.e., $\mathcal{D}_{i}=\mathcal{D}_{j}$ for all $i,j\in[N]$. In this work, we focus on the more challenging heterogeneous (non-i.i.d.) case: i.e. the data distributions $\{\mathcal{D}_{i}\}_{i=1}^{N}$ may be arbitrary. This case arises in many practical FL settings Kairouz et al. (2021).
@@
The quality of a private algorithm $\mathcal{A}$ for solving Problem FL is measured by its excess risk (a.k.a. excess population/test loss), $\mathbb{E}[F(\mathcal{A}(\mathbf{X}))]-F^{*}$, where $F^{*}=\inf_{w\in\mathcal{W}}F(w)$ and the expectation is over the random draw of the data $\mathbf{X}=(X_{1},\ldots,X_{N})$ as well as the randomness of $\mathcal{A}$Lowy and Razaviyayn (2023a) characterized the minimax optimal excess risk (up to logarithms) for convex ISRL-DP FL in the homogeneous (i.i.d.) case:
@@
where $d$ is the dimension of the parameter space and $\varepsilon,\delta$ are the privacy parameters. For heterogeneous FL, the state-of-the-art excess risk bound is $O\left(1/\sqrt{Nn}+\left(\sqrt{d\log(1/\delta)}/(\varepsilon n\sqrt{N})\right)^{4/5}\right)$, assuming smoothness of the loss function Lowy and Razaviyayn (2023a). Lowy and Razaviyayn (2023a) left the following question open:
@@
In practical FL settings (especially cross-device settings), it can be expensive or slow for silos to communicate with the server or with each other Kairouz et al. (2021). Thus, another important desiderata for FL algorithms is communication efficiency, which is measured by communication complexity: the number of communication rounds $R$ that are required to achieve excess risk $\alpha$. In the homogeneous setting, the state-of-the-art communication complexity for an algorithm achieving optimal excess risk is due to Lowy and Razaviyayn (2023a, Theorem D.1):
@@
When $N=1$ in (FL), ISRL-DP FL reduces to central DP stochastic convex optimization (SCO), which has been studied extensively Bassily et al. (2019); Feldman et al. (2020); Asi et al. (2021); Zhang et al. (2022). In this centralized setting, communication complexity is usually referred to as iteration complexity. Even in the special case of $N=1$, our iteration complexity for smooth losses improves over the prior state-of-the-art result Zhang et al. (2022).
@@
In the homogeneous case, the state-of-the-art gradient complexity for an ISRL-DP FL algorithm that attains optimal excess risk is due to Lowy and Razaviyayn (2023a, Theorem D.1). For smooth losses and $\varepsilon=\Theta(1)$, Lowy et al. (2023a) obtained the follow gradient complexity
@@
In Theorem 3.4, we also improve over the state-of-the-art subgradient complexity bound of Lowy and Razaviyayn (2023a, Theorem D.2) for non-smooth losses. Moreover, in contrast to these earlier results, we do not assume homogeneous data.
@@
Table 1: Comparison vs. SOTA for Smooth Loss Functions. In Tables 1-2, [LR’23] refers to Lowy and Razaviyayn (2023a); we omit logs and fix $M=N$, $d=\Theta(n)$, $\varepsilon=\Theta(1)$, $L=\beta=D=1$. See theorems and the appendix for more general cases.
@@
Our ISRL-DP Algorithm 1 for smooth FL builds on the ISRL-DP Accelerated MB-SGD approach used by Lowy and Razaviyayn (2023a) for federated empirical risk minimization (ERM). We call this algorithm repeatedly to iteratively solve a carefully chosen sequence of regularized ERM subproblems, using the localization technique of Feldman et al. (2020). We obtain our novel optimal excess risk bounds for heterogeneous FL with an algorithmic stability Bousquet and Elisseeff (2002) argument. We make the key observation that the stability and generalization of regularized ERM in Shalev-Shwartz et al. (2009) does not require the data to be identically distributed. By combining this observation with a bound for ISRL-DP Accelerated MB-SGD, we obtain our excess risk bound.
It is not immediately clear that the approach just described should work, because the iterates of Accelerated MB-SGD are not necessarily stable. The instability of acceleration may explain why Lowy and Razaviyayn (2023a) used ISRL-DP Accelerated MB-SGD only for ERM and not for minimizing the population risk. We overcome this obstacle with an alternative analysis that leverages the stability of regularized ERM, the convergence of ISRL-DP Accelerated MB-SGD to an approximate minimizer of the regularized empirical loss, and localization. This argument enables us to show that our algorithm has optimal excess risk. By carefully choosing algorithmic parameters, we also obtain state-of-the-art communication and gradient complexities.
To extend our algorithm to the non-smooth case, we use two different techniques. One is Nesterov smoothing Nesterov (2005), which results in an algorithm with favorable communication complexity. The other approach is to replace the ISRL-DP Accelerated MB-SGD ERM subsolver by an ISRL-DP Subgradient Method. This technique yields an algorithm with favorable (sub)gradient complexity. Third, we use randomized convolution smoothing, as in Kulkarni et al. (2021), to obtain another favorable gradient complexity bound.
@@ 1.1 Preliminaries
Definition 1.1 (Differential Privacy Dwork et al. (2006)).
@@
We assume the following throughout:
@@ 2 Localized ISRL-DP Accelerated MB-SGD for Smooth Losses
We start with the smooth case. Combining iterative localization techniques of Feldman et al. (2020); Asi et al. (2021) with the multi-stage ISRL-DP Accelerated MB-SGD of Lowy and Razaviyayn (2023a), our proposed Algorithm 1 achieves optimal excess risk and state-of-the-art communication complexity for heterogeneous FL under ISRL-DP.
@@
Algorithm 2 Accelerated ISRL-DP MB-SGD Lowy and Razaviyayn (2023a)
1:Datasets $X_{l}\in\mathcal{X}^{n}$ for $l\in[N]$, loss function $\hat{F}(w)=\frac{1}{nN}\sum_{l=1}^{N}\sum_{x\in X_{l}}f(w,x)$, constraint set $\mathcal{W}$, initial point $w_{0}$, strong convexity modulus $\mu\geq 0$, privacy parameteres $(\varepsilon,\delta)$, iteration count $R\in\mathbb{N}$, batch size $K\in[n]$, step size parameters $\{\eta_{r}\}_{r\in[R]},\{\alpha_{r}\}_{r\in[R]}$ specified in Section 7.
@@
We recall ISRL-DP Accelerated MB-SGD in Algorithm 2. It is a distributed, ISRL-DP version of the AC-SA algorithm Ghadimi and Lan (2012). For strongly convex losses, the multi-stage implementation of ISRL-DP Accelerated MB-SGD, given in Algorithm 5 in Section 7, offers improved excess risk Lowy and Razaviyayn (2023a). Algorithm 5 is a distributed, ISRL-DP version of the multi-stage AC-SA Ghadimi and Lan (2013).
Building on Algorithm 5, we describe our algorithm as follows (see Algorithm 1 for pseudocode). The distributed learning process is divided into $\tau=\lfloor\log_{2}n\rfloor$ phases. In each phase $i\in[\tau]$, all silos work together (via communication with the central server) to iteratively solve a regularized ERM problem with ISRL-DP. The regularized ERM problem in phase $i$ is defined over $N$ local batches of data, each containing $n_{i}$ disjoint samples (cf. $\hat{F}_{i}(w)$ in Line 7). To find the approximate constrained minimizer of $\hat{F}_{i}(w)$ privately, we apply Algorithm 5 for a careful choice of the number of rounds $R_{i}$ and the batch size $K_{i}\in[n_{i}]$ (cf. Line 8). The output $w_{i}$ of phase $i$ affects phase $i+1$ in three ways: (i) regularization center used to define $\hat{F}_{i+1}$; (ii) initialization for the next call of Algorithm 5; and (iii) constraint set $\mathcal{W}_{i+1}$. We enforce stability (hence generalization) of our algorithm via regularization and localization: e.g., as $i$ increases, we increase the regularization parameter $\lambda_{i}$ to prevent $w_{i+1}$ from moving too far away from $w_{i}$ and $w^{*}$.
@@
For general $d,n,\varepsilon$, the gradient complexity expression is complicated, and is given in the Appendix 8.1.
@@
The excess risk bound in (9) matches the optimal i.i.d risk bound up to log factors (cf. Theorem 2.2 in Lowy and Razaviyayn (2023a)), even when silo data is arbitrarily heterogeneous across silos. To the best of our knowledge, our algorithm is the first to have this property, resolving an open question of Lowy and Razaviyayn (2023a). The prior state-of-the-art bound in Lowy and Razaviyayn (2023a, Theorem 3.1) is suboptimal by a factor of $\widetilde{O}((\sqrt{d}/(\varepsilon n\sqrt{N}))^{1/5})$.
@@
Our algorithm has state-of-the-art communication complexity, even in the simple case of $N=1$, where ISRL-DP FL reduces to central DP SCO. In fact, the prior state-of-the-art iteration complexity bound for DP SCO was $O((\beta D/L)\min\{\sqrt{n},\varepsilon n/\sqrt{d\ln(1/\delta)}\})$ Zhang et al. (2022). By comparison, when $N=1$, our algorithm’s communication complexity is the square root of this bound. Note that when $N=1$, our algorithm is essentially the same as the algorithm of Kulkarni et al. (2021), but we do not incorporate convolution smoothing here since we are assuming smoothness of the loss.
We can also compare our gradient complexity results against the state-of-the-art central DP SCO algorithms when $N=1$. As an illustration, consider the interesting regime $\varepsilon=\Theta(1)$ and $d=\Theta(n)$. For smooth and convex losses, when (i) $\beta\leq O(\sqrt{n}L/D)$, algorithms in both Feldman et al. (2020) and Zhang et al. (2022) achieve optimal risk using $\widetilde{O}(n)$ gradient complexity. For (ii) $\beta\geq\Omega(\sqrt{n}L/D)$, the two algorithms proposed by Feldman et al. (2020) fail to guarantee optimal risk, whereas (Zhang et al., 2022, Algorithm 2) continues to attain optimal risk with gradient complexity $\widetilde{O}(n^{3/4}\sqrt{\beta D/L})$. By comparison with these results, for case (i), our Algorithm 1 achieves optimal risk with gradient complexity $\widetilde{O}(n^{5/4})$. For case (ii), as in (Zhang et al., 2022), our Algorithm 1 continues to achieve optimal risk with gradient complexity $\widetilde{O}(n^{9/8}(\beta D/L)^{1/4}+n^{1/4}(\beta D/L)^{1/2}+n)$. Thus, our algorithm is faster than the state-of-the-art result of Zhang et al. (2022) when $\beta D/L\geq n^{3/2}$. In the complementary parameter regime, however, the algorithm of Zhang et al. (2022) (which is not ISRL-DP) is faster. We discuss possible ways to close this gap in Section 5.
@@
Theorem 2.1 trivially extends to the unconstrained optimization setting in which $D=\|w_{0}-w^{*}\|$ for $w^{*}\in\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}F(w)$. Moreover, our excess risk bound is still optimal for the unconstrained case: a matching lower bound is obtained by combining the technique for unconstrained DP lower bounds in Liu and Lu (2021) with the constrained ISRL-DP FL lower bounds of Lowy and Razaviyayn (2023a).
@@
The lower bound holds for the class of distributed zero-respecting algorithms (defined in Appendix 11), which includes most non-private FL algorithms, such as MB-SGD, Accelerated MB-SGD, local SGD/FedAvg, and so on.
Theorem 2.4 (Communication Lower Bound Woodworth et al. (2020)).
@@
where $\hat{w}_{i}=\operatorname*{arg\,min}_{w\in\mathcal{W}}\hat{F}_{i}(w)$ for $i\in[\tau]$. Then, it remains to bound (10) and (11), respectively. To this end, we first show that $w_{i}$ in Algorithm 1 is close to $\hat{w}_{i}$ for all $i\in[\tau]$. Here, we mainly leverage the excess empirical risk bound of Algorithm 5 for ERM with strongly convex and smooth loss functions. In particular, by the empirical risk bound in Lowy and Razaviyayn (2023a, Theorem F.1) and $\lambda_{i}$-strongly convex of $\hat{F}_{i}$, we can show that the following bound holds for all $i\in[\tau]$:
@@
We obtain a bounds on (11) by combining the above inequality with (12). Finally, by putting everything together and leveraging the geometrical schedule of $\lambda_{i},n_{i}$, we obtain the final excess population risk bound. ∎
An alternative proof approach would be to try to show that Algorithm 5 is stable directly, e.g. by using the tools that Hardt et al. (2016) use for proving stability of SGD. However, this approach seems unlikely to yield the same tight bounds that we obtain, since acceleration may impede stability of the iterates. Instead, we establish our stability and generalization guarantee indirectly: We use the facts that reguralized ERM is stable and that Algorithm 5 approximates regularized ERM.

### v2

@@ Abstract
We revisit the problem of federated learning (FL) with private data from people who do not trust the server or other silos/clients. In this context, every silo (e.g. hospital) has data from several people (e.g. patients) and needs to protect the privacy of each person’s data (e.g. health records), even if the server and/or other silos try to uncover this data. Inter-Silo Record-Level Differential Privacy (ISRL-DP) prevents each silo’s data from being leaked, by requiring that silo $i$’s communications satisfy item-level differential privacy. Prior work (Lowy and Razaviyayn, 2023b) characterized the optimal excess risk bounds for ISRL-DP algorithms with homogeneous (i.i.d.) silo data and convex loss functions. However, two important questions were left open: (1) Can the same excess risk bounds be achieved with heterogeneous (non-i.i.d.) silo data? (2) Can the optimal risk bounds be achieved with fewer communication rounds? In this paper, we give positive answers to both questions. We provide novel ISRL-DP FL algorithms that achieve the optimal excess risk bounds in the presence of heterogeneous silo data. Moreover, our algorithms are more communication-efficient than the prior state-of-the-art. For smooth loss functions, our algorithm achieves the optimal excess risk bound and has communication complexity that matches the non-private lower bound. Additionally, our algorithms are more computationally efficient than the previous state-of-the-art.
@@ 1 Introduction
Federated learning (FL) is a distributed machine learning paradigm in which multiple silos (a.k.a. clients), such as hospitals or cell-phone users, collaborate to train a global model. In FL, silos store their data locally and exchange focused updates (e.g. stochastic gradients), sometimes making use of a central server (Kairouz et al., 2021). FL has been applied in myriad domains, from consumer digital products (e.g. Google’s mobile keyboard (Hard et al., 2018) and Apple’s iOS (Apple, 2019)) to healthcare (Courtiol et al., 2019), finance (FedAI, 2019), and large language models (LLMs) (Hilmkil et al., 2021).
@@
Differential privacy (DP) (Dwork et al., 2006) guarantees that private data cannot be leaked. Different variations of DP have been considered for FL. Central DP prevents the final trained FL model from leaking data to an external adversary (Jayaraman and Wang, 2018; Noble et al., 2022). However, central DP has two major drawbacks: (1) it does not provide a privacy guarantee for each individual silo; and (2) it does not ensure privacy when an attacker/eavesdropper has access to the server or to another silo.
Another notion of DP for FL is user-level DP (McMahan et al., 2018; Geyer et al., 2017; Levy et al., 2021). User-level DP mitigates the drawback (1) of central DP, by providing privacy for the complete local data set of each silo. User-level DP is practical in the cross-device FL setting, where each silo is a single person (e.g. cell-phone user) with a large number of records (e.g. text messages). However, user-level DP still permits privacy breaches if an adversary has access to the server or eavesdrops on the communications between silos. Moreover, user-level DP is not well-suited for cross-silo federated learning, where silos represent organizations like hospitals, banks, or schools that house data from many individuals (e.g. patients, customers, or students). In the cross-silo FL context, each person possesses a record, referred to as an “item,” which may include sensitive data. Therefore, an appropriate notion of DP for cross-silo FL should safeguard the privacy of each individual record (i.e. “item-level differential privacy”) within silo $i$, rather than the complete aggregated data of silo $i$.
Following Lowy et al. (2023a); Lowy and Razaviyayn (2023b); Liu et al. (2022) (among others), this work considers inter-silo record-level DP (ISRL-DP). ISRL-DP requires that the full transcript of messages sent by silo $i$ satisfy item-level DP, for all silos $i$. Thus, ISRL-DP guarantees the privacy of each silo’s local data, even in the presence of an adversary with access to the server, other silos, or the communication channels between silos and the server. See Figure 1 for an illustration. If each silo only has one record, then ISRL-DP reduces to local DP (Kasiviswanathan et al., 2011; Duchi et al., 2013). However, in the FL setting where each silo has many records, ISRL-DP FL is more appropriate and permits much higher accuracy than local DP (Lowy and Razaviyayn, 2023b). Moreover, ISRL-DP implies user-level DP if the ISRL-DP parameters are sufficiently small (Lowy and Razaviyayn, 2023b). Thus, ISRL-DP is a practical privacy notion for cross-silo and cross-device FL when individuals do not trust the server or other silos/clients/devices with their sensitive data.
@@
Assume that the samples $\{x_{i,j}\}_{i\in[N],j\in[n]}$ are independent. The FL problem is homogeneous if the data is i.i.d., i.e., $\mathcal{D}_{i}=\mathcal{D}_{j}$ for all $i,j\in[N]$. In this work, we focus on the more challenging heterogeneous (non-i.i.d.) case: i.e. the data distributions $\{\mathcal{D}_{i}\}_{i=1}^{N}$ may be arbitrary. This case arises in many practical FL settings (Kairouz et al., 2021).
@@
The quality of a private algorithm $\mathcal{A}$ for solving Problem FL is measured by its excess risk (a.k.a. excess population/test loss), $\mathbb{E}[F(\mathcal{A}(\mathbf{X}))]-F^{*}$, where $F^{*}=\inf_{w\in\mathcal{W}}F(w)$ and the expectation is over the random draw of the data $\mathbf{X}=(X_{1},\ldots,X_{N})$ as well as the randomness of $\mathcal{A}$Lowy and Razaviyayn (2023b) characterized the minimax optimal excess risk (up to logarithms) for convex ISRL-DP FL in the homogeneous (i.i.d.) case:
@@
where $d$ is the dimension of the parameter space and $\varepsilon,\delta$ are the privacy parameters. For heterogeneous FL, the state-of-the-art excess risk bound is $O\left(1/\sqrt{Nn}+\left(\sqrt{d\log(1/\delta)}/(\varepsilon n\sqrt{N})\right)^{4/5}\right)$, assuming smoothness of the loss function (Lowy and Razaviyayn, 2023b). Lowy and Razaviyayn (2023b) left the following question open:
@@
In practical FL settings (especially cross-device settings), it can be expensive or slow for silos to communicate with the server or with each other (Kairouz et al., 2021). Thus, another important desiderata for FL algorithms is communication efficiency, which is measured by communication complexity: the number of communication rounds $R$ that are required to achieve excess risk $\alpha$. In the homogeneous setting, the state-of-the-art communication complexity for an algorithm achieving optimal excess risk is due to Lowy and Razaviyayn (2023b, Theorem D.1):
@@
When $N=1$ in (FL), ISRL-DP FL reduces to central DP stochastic convex optimization (SCO), which has been studied extensively (Bassily et al., 2019; Feldman et al., 2020; Asi et al., 2021; Zhang et al., 2022). In this centralized setting, communication complexity is usually referred to as iteration complexity. Even in the special case of $N=1$, our iteration complexity for smooth losses improves over the prior state-of-the-art result (Zhang et al., 2022).
@@
In the homogeneous case, the state-of-the-art gradient complexity for an ISRL-DP FL algorithm that attains optimal excess risk is due to Lowy and Razaviyayn (2023b, Theorem D.1). For smooth losses and $\varepsilon=\Theta(1)$, Lowy and Razaviyayn (2023b) obtained the follow gradient complexity
@@
In Theorem 3.4, we also improve over the state-of-the-art subgradient complexity bound of Lowy and Razaviyayn (2023b, Theorem D.2) for non-smooth losses. Moreover, in contrast to these earlier results, we do not assume homogeneous data.
@@
Table 1: Comparison vs. SOTA for Smooth Loss Functions. [LR’23] refers to Lowy and Razaviyayn (2023b); we omit logs and fix $M=N$, $d=\Theta(n)$, $\varepsilon=\Theta(1)$, $L=\beta=D=1$. See theorems and the appendix for more general cases.
@@
Our ISRL-DP Algorithm 1 for smooth FL builds on the ISRL-DP Accelerated MB-SGD approach used by Lowy and Razaviyayn (2023b) for federated empirical risk minimization (ERM). We call this algorithm repeatedly to iteratively solve a carefully chosen sequence of regularized ERM subproblems, using the localization technique of Feldman et al. (2020). We obtain our novel optimal excess risk bounds for heterogeneous FL with an algorithmic stability (Bousquet and Elisseeff, 2002) argument. The key observation is that the stability and generalization of regularized ERM in Shalev-Shwartz et al. (2009) does not require the data to be identically distributed. By combining this observation with a bound for ISRL-DP Accelerated MB-SGD, we obtain our excess risk bound.
It is not immediately clear that the approach just described should work, because the iterates of Accelerated MB-SGD are not necessarily stable. The instability of acceleration may explain why Lowy and Razaviyayn (2023b) used ISRL-DP Accelerated MB-SGD only for ERM and not for minimizing the population risk. We overcome this obstacle with an alternative analysis that leverages the stability of regularized ERM, the convergence of ISRL-DP Accelerated MB-SGD to an approximate minimizer of the regularized empirical loss, and localization. This argument enables us to show that our algorithm has optimal excess risk. By carefully choosing algorithmic parameters, we also obtain state-of-the-art communication and gradient complexities.
To extend our algorithm to the non-smooth case, we use two different techniques. One is Nesterov smoothing (Nesterov, 2005), which results in an algorithm with favorable communication complexity. The other approach is to replace the ISRL-DP Accelerated MB-SGD ERM subsolver by an ISRL-DP Subgradient Method. This technique yields an algorithm with favorable (sub)gradient complexity. Third, we use randomized convolution smoothing, as in Kulkarni et al. (2021), to obtain another favorable gradient complexity bound.
@@ 1.1 Preliminaries
Definition 1.1 (Differential Privacy (Dwork et al., 2006)).
@@
Assume the following throughout:
@@ 2 Localized ISRL-DP Accelerated MB-SGD for Smooth Losses
We start with the smooth case. Combining iterative localization techniques of Feldman et al. (2020); Asi et al. (2021) with the multi-stage ISRL-DP Accelerated MB-SGD of Lowy and Razaviyayn (2023b), our proposed Algorithm 1 achieves optimal excess risk and state-of-the-art communication complexity for heterogeneous FL under ISRL-DP.
@@
Algorithm 2 Accelerated ISRL-DP MB-SGD (Lowy and Razaviyayn, 2023b)
1:Datasets $X_{l}\in\mathcal{X}^{n}$ for $l\in[N]$, loss function $\hat{F}(w)=\frac{1}{nN}\sum_{l=1}^{N}\sum_{x\in X_{l}}f(w,x)$, constraint set $\mathcal{W}$, initial point $w_{0}$, strong convexity modulus $\mu\geq 0$, privacy parameters $(\varepsilon,\delta)$, iteration count $R\in\mathbb{N}$, batch size $K\in[n]$, step size parameters $\{\eta_{r}\}_{r\in[R]},\{\alpha_{r}\}_{r\in[R]}$ specified in Appendix B.
@@
We recall ISRL-DP Accelerated MB-SGD in Algorithm 2. It is a distributed, ISRL-DP version of the AC-SA algorithm (Ghadimi and Lan, 2012). For strongly convex losses, the multi-stage implementation of ISRL-DP Accelerated MB-SGD, given in Algorithm 5 in Appendix B, offers improved excess risk (Lowy and Razaviyayn, 2023b). Algorithm 5 is a distributed, ISRL-DP version of the multi-stage AC-SA (Ghadimi and Lan, 2013).
Building on Algorithm 5, we describe our algorithm as follows (see Algorithm 1 for pseudocode). The distributed learning process is divided into $\tau=\lfloor\log_{2}n\rfloor$ phases. In each phase $i\in[\tau]$, all silos work together (via communication with the central server) to iteratively solve a regularized ERM problem with ISRL-DP. The regularized ERM problem in phase $i$ is defined over $N$ local batches of data, each containing $n_{i}$ disjoint samples (cf. $\hat{F}_{i}(w)$ in 7). To find the approximate constrained minimizer of $\hat{F}_{i}(w)$ privately, we apply Algorithm 5 for a careful choice of the number of rounds $R_{i}$ and the batch size $K_{i}\in[n_{i}]$ (cf. 8). The output $w_{i}$ of phase $i$ affects phase $i+1$ in three ways: (i) regularization center used to define $\hat{F}_{i+1}$; (ii) initialization for the next call of Algorithm 5; and (iii) constraint set $\mathcal{W}_{i+1}$. We enforce stability (hence generalization) of our algorithm via regularization and localization: e.g., as $i$ increases, we increase the regularization parameter $\lambda_{i}$ to prevent $w_{i+1}$ from moving too far away from $w_{i}$ and $w^{*}$.
@@
For general $d,n,\varepsilon$, the gradient complexity expression is complicated, and is given in the Section C.1.
@@
The excess risk bound in (9) matches the optimal i.i.d risk bound up to log factors (cf. Theorem 2.2 in Lowy and Razaviyayn (2023b)), even when silo data is arbitrarily heterogeneous across silos. To the best of our knowledge, our algorithm is the first to have this property, resolving an open question of Lowy and Razaviyayn (2023b). The prior state-of-the-art bound in Lowy and Razaviyayn (2023b, Theorem 3.1) is suboptimal by a factor of $\widetilde{O}((\sqrt{d}/(\varepsilon n\sqrt{N}))^{1/5})$.
@@
Our algorithm has state-of-the-art communication complexity, even in the simple case of $N=1$, where ISRL-DP FL reduces to central DP SCO. In fact, the prior state-of-the-art iteration complexity bound for DP SCO was $O((\beta D/L)\min\{\sqrt{n},\varepsilon n/\sqrt{d\ln(1/\delta)}\})$ (Zhang et al., 2022). By comparison, when $N=1$, our algorithm’s communication complexity is the square root of this bound. Note that when $N=1$, our algorithm is essentially the same as the algorithm of Kulkarni et al. (2021), but we do not incorporate convolution smoothing here since we are assuming smoothness of the loss.
We can also compare our gradient complexity results against the state-of-the-art central DP SCO algorithms when $N=1$. As an illustration, consider the interesting regime $\varepsilon=\Theta(1)$ and $d=\Theta(n)$. For smooth and convex losses, when (i) $\beta\leq O(\sqrt{n}L/D)$, algorithms in both Feldman et al. (2020) and Zhang et al. (2022) achieve optimal risk using $\widetilde{O}(n)$ gradient complexity. For (ii) $\beta\geq\Omega(\sqrt{n}L/D)$, the two algorithms proposed by Feldman et al. (2020) fail to guarantee optimal risk, whereas Zhang et al. (2022, Algorithm 2) continues to attain optimal risk with gradient complexity $\widetilde{O}(n^{3/4}\sqrt{\beta D/L})$. By comparison with these results, for case (i), our Algorithm 1 achieves optimal risk with gradient complexity $\widetilde{O}(n^{5/4})$. For case (ii), as in Zhang et al. (2022), our Algorithm 1 continues to achieve optimal risk with gradient complexity $\widetilde{O}(n^{9/8}(\beta D/L)^{1/4}+n^{1/4}(\beta D/L)^{1/2}+n)$. Thus, our algorithm is faster than the state-of-the-art result of Zhang et al. (2022) when $\beta D/L\geq n^{3/2}$. In the complementary parameter regime, however, the algorithm of Zhang et al. (2022) (which is not ISRL-DP) is faster. We discuss possible ways to close this gap in Section 5.
@@
Theorem 2.1 trivially extends to the unconstrained optimization setting in which $D=\|w_{0}-w^{*}\|$ for $w^{*}\in\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}F(w)$. Moreover, our excess risk bound is still optimal for the unconstrained case: a matching lower bound is obtained by combining the technique for unconstrained DP lower bounds in Liu and Lu (2021) with the constrained ISRL-DP FL lower bounds of Lowy and Razaviyayn (2023b).
@@
The lower bound holds for the class of distributed zero-respecting algorithms (defined in Appendix F), which includes most non-private FL algorithms, such as MB-SGD, Accelerated MB-SGD, local SGD/FedAvg, and so on.
Theorem 2.4 (Communication Lower Bound (Woodworth et al., 2020)).
@@
where $\hat{w}_{i}=\operatorname*{arg\,min}_{w\in\mathcal{W}}\hat{F}_{i}(w)$ for $i\in[\tau]$. Then, it remains to bound (10) and (11), respectively. To this end, we first show that $w_{i}$ in Algorithm 1 is close to $\hat{w}_{i}$ for all $i\in[\tau]$. Here, we mainly leverage the excess empirical risk bound of Algorithm 5 for ERM with strongly convex and smooth loss functions. In particular, by the empirical risk bound in Lowy and Razaviyayn (2023b, Theorem F.1) and $\lambda_{i}$-strongly convex of $\hat{F}_{i}$, we can show that the following bound holds for all $i\in[\tau]$:
@@
We obtain a bound on (11) by combining the above inequality with (12). Finally, by putting everything together and leveraging the geometrical schedule of $\lambda_{i},n_{i}$, we obtain the final excess population risk bound. ∎
An alternative proof approach would be to try to show that Algorithm 5 is stable directly, e.g. by using the tools that Hardt et al. (2016) use for proving stability of SGD. However, this approach seems unlikely to yield the same tight bounds that we obtain, since acceleration may impede stability of the iterates. Instead, we establish our stability and generalization guarantee indirectly: We use the facts that regularized ERM is stable and that Algorithm 5 approximates regularized ERM.